Suppose the IS curve for Fiji is Y = 2500-125i and the LM is Y = 1000+25i. Use Cramer’s rule to determine i* and Y*. Now suppose the Fijian Government increased spending by $100m. Determine the extent of crowding out of private investment and the increase in Ms required to dampen any crowding-out effect. You are given the following parameters: αG = 2.5, h = 65, k = 0.5, b = 50.
1
Expert's answer
2020-06-11T11:18:40-0400

"1) Y = 1250 , i = 10%"

"2) 233.5"

3) Increase in money supply = 100

Explanation:

The equation for IS:

"Y = 2500-125i ......(1)"

The equation for LM:

"Y = 1000+25i .......(2)"

Other given information are:

"\u03b1G = 2.5, \nh = 65, \nk = 0.5, \nb = 50."

From equation;"(1): Y + 125i = 2500"

From equation;(2): Y - 25i = 1000

where:a1 = 1 , a2 = 1, b1 = 125 , b2 = -25 , c1 = 2500 , c2 = 1000


From Cramers rule 5:

"Y = | [2500(-25) - 1000(125)] \/ [1(-25) - 1(125)] |"

"= | (-62500 - 125000) \/ (-25 - 125) |"

"= 187500\/150"

"= 1250"

By substituting the value of Y in equation (2):

"i = 10%"


2)

change in Y = government spending multiplier * change in G

"= \u03b1G *100"

"= 2.5 * 100"

"= 250"

It implies that due to an increase in G by $100m , the increase in Y should be 250.


But due to the crowd out of private investment, the increase in Y is less. The real increase in Y can be calculated as follows:

New equation of IS:

"Y = 2500 - 125i + 100"

"Y = 2600 - 125i"

The equation of LM is same as before:

"Y = 1000+25i"

By solving these two equations:

"Y = 1266.5 and\ni = 10.66%"

The real change in Y = 1266.5 - 1250 = 16.5 

Crowd out "= 250 - 16.5 = 233.5"


3)

The equation of Lm explains the money market equilibrium:

Money supply = money demand 

"MS = kY - hi"

Suppose with the increase in money supply to offset the crowd out effect, the new money supply is MS'.

"MS' =kY - hi"

"MS' = 0.5Y - 65i"

At the new equilibrium, we want an increase in Y of 250 and i "= 10%"

"MS' = 0.5*(1250 + 250) - 65*(10)"

"MS' = 0.5(1500) - 650"

"MS' = 750 - 650"

"MS' = 100"


Thus, an increase in money supply should be 100.